OpenStudy (anonymous):

Can somebody help me with this problem: If y^3=25x^2, determine dx/dt when x=5 and dy/dt=-2

OpenStudy (anonymous):

Implicit differentiation: 3y^2 dy/dt = 50x dx/dt Now plug in........

OpenStudy (anonymous):

3y^2(-2)=50(-5)dx/dt

OpenStudy (anonymous):

We don't know 'y'

OpenStudy (anonymous):

You are right...but you do know that x = 5, so substitute x = 5 into the original equation, and solve for y (that's part of the problem).

OpenStudy (anonymous):

ah!

OpenStudy (anonymous):

you get y^3 = 625 y = cube root of 625 or (625)^(1/3)

OpenStudy (anonymous):

Now, you can easily finish to find dx/dt

OpenStudy (anonymous):

Thanks for the help.

OpenStudy (anonymous):

welcome.